OpenStudy (anonymous):

How do I solve 8^2x=2^x+5?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

write \(8=2^3\) s o \[\huge 8^{2x}=(2^3)^{2x}=2^{6x}\]

OpenStudy (misty1212):

then if \[2^{6x}=2^{x+5}\] you know ' \[\huge 6x=x+5\]

OpenStudy (usukidoll):

what the O_O

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (misty1212):

\[\huge \color\magenta \heartsuit\]